# A Study of Undefined Behavior Across Foreign Function Boundaries in Rust Libraries

### v4

@@ IV-B Typing
We found bugs from crates involving values which were partially initialized or had invalid alignment. Most were caused by incorrect use of MaybeUninit<T>, but several involved incorrect foreign function bindings.
@@
We found bugs involving partial initialization of MaybeUninit<T>. Bugs #0 and #0 were each found in the same library. In each case, an uninitialized instance of a struct T containing an array of values was created using MaybeUninit<T> and passed across the FFI to be initialized. The foreign call only initialized the first element of the array to 0. The remaining uninitialized elements would never be read by C, since every iteration stopped at the null terminator. Miri reported these initialization patterns as an error when MaybeUninit<T>::assume_init() was called. We fixed each bug by zero-initializing the entire array in Rust.
Bug #0 seemed to involve an equivalent pattern; MaybeUninit<T>::assume_init() triggered an error for a struct that was initialized by a foreign function. However, the function properly zero-initialized the struct using memset. After examining the LLVM bytecode of the foreign function, we found that all 88 bytes of the struct were initialized on the LLVM stack, but only the first 80 bytes were copied back into Rust. These missing bytes corresponded to padding fields, which were never directly accessed by the foreign library. Optimizations had been disabled, but LLVM still removed this ‘‘unnecessary’’ write access. We fixed Bug #0 by zero-initializing the padding fields before calling assume_init().
Incorrect FFI Bindings
We found crates with one or more incorrect foreign function bindings. All bindings from each crate had been written manually. Three of these crates had bindings with missing return types. In Bug #0, a binding was declared in a unit test without its 32-bit integer return type. The C implementation used this to return a status code indicating if one of its integer parameters was within bounds. The remaining two bugs involved bindings with incorrect integer types, which lead to incomplete initialization. In Bug #0, a function was declared to return a 32-bit integer, but its C implementation returned a 1-byte boolean value. In Bug #0, the last parameter of a function was declared as a 32-bit integer, but the C implementation expected a size_t. The width of size_t is architecture dependent, so this binding was only correct for 32-bit architectures. To fix this bug, multiple prior releases needed to be patched to use Rust’s usize.
@@ IV-C Allocation
We found issues related to allocation in distinct crates. These include memory leaks and 1 new cross-language deallocation bug. We also found null-pointer dereference, but it did not have any connection to Rust.
@@
Rust’s Box<U> and CString use similar APIs to encapsulate heap allocations. A new instance is created with T::new and deallocated when it leaves scope. The function T::into_raw consumes an instance of T, producing a pointer to the interior allocation. To avoid a memory leak, this pointer must be reconstituted in its wrapper type using T::from_raw. We found leaks caused by calling into_raw on either Box<T> and CString without later calling from_raw. Each type was used to allocate memory for a foreign function call, and each bug was fixed by adding a call to T::from_raw after the foreign function returned. We also encountered bugs involving leaks of memory allocated by C. Each was caused by neglecting to call the appropriate destructor function exposed by the C API. However, in Bug #0, the C API did not expose a destructor since it had been designed under the assumption that users would be able to call free. We could not use Rust’s std::alloc::dealloc to fix this, since that could lead to invalid cross-language deallocation.
@@ V Discussion
Our findings indicate that it is easy to inadvertently introduce undefined behavior in Rust libraries that call foreign functions. The errors we found can be prevented with careful auditing, but it is seems likely that many will persist, undetected, until the Rust community develops a production-ready method for finding aliasing violations in multi-language applications. We provide the following recommendations:
@@
The Rust community is in dire need of a production-ready solution for validating Rust’s aliasing model across foreign function boundaries. Though our approach was capable of finding useful results, its scalability is limited due to the requirement of implementing shims for system calls and the lack of a formal specification for the ABIs implemented by Rust, Clang, and LLVM. Instrumenting a shared, intermediate format is likely to be the most effective approach. The Krabcake project [33] is developing an extension to Valgrind [34] (concurrent to MiriLLI) to provide support for detecting aliasing violations, but it is not yet capable of replicating our results. The Rust community should invest resources into completing a prototype implementation to evaluate on real-world applications. However, since Valgrind can incur significant runtime overhead, it would also be worthwhile to develop an equivalent approach using LLVM’s Sanitizer API [35].
@@ VI Related Work
Studies that examined unsafe code in Rust libraries have consistently found that foreign function calls are a significant use-case for unsafe code [4, 5]. In particular, Evans et al. [4] surveyed all published crates in September, 2018 and found that 22.5% of all unsafe function calls were to foreign functions. Rust developers also view foreign function calls as a central use case for unsafe code. Fulton et al [6] interviewed 16 Rust developers and surveyed 178 developers, finding that nearly half of participants interoperated with foreign libraries. Both Höltervennhoff et al. [7] and McCormack et al. [8] focused specifically on Rust developers who use unsafe code, and each found that the majority of participants used Rust’s FFI.
@@
Prior studies have examined open-source contributions to categorize errors caused by unsafe code. Most considered foreign function calls to be out-of-scope, and none described aliasing violations in terms of Stacked or Tree Borrows. Both Qin et al. [36] and Xu et al. [37] examined bug and vulnerability reports from open source projects and found examples of allocation and typing errors similar to the ones that we identified. Xu et al. [37] did encounter errors related to FFI use, but they were primarily ‘‘straightforward’’ issues related to application-specific invariants, layout, and alignment. Cui et al. [38] evaluated a taxonomy of 19 safety properties required by unsafe functions, but they made it an explicit goal to “disregard” foreign functions.
@@
Prior approaches to static analysis [39, 36, 40, 41] have found a wide variety of bugs caused by improper use of unsafe code, but few approaches scale to multi-language applications. Li et al. [42] applied dataflow analysis on LLVM IR and detected several memory leaks and cross-language deallocation errors in Rust libraries. We replicated several of the bugs that they encountered. Hu et al. [43] modified several existing Rust analysis tools, including Miri, to analyze multi-language programs defined in a custom intermediate representation. However, they did not evaluate their approach on any real Rust libraries and did not consider aliasing violations. Lei et al. [44] took a similar approach using WebAssembly as a target language; their evaluation did not use Miri or consider aliasing issues. Recent deductive verifiers approaches based on the RustBelt [45] model have used compositional verification [46] and automated proof search [47] to verify unsafe code, but no static verifiers reason about Stacked or Tree Borrows.
@@ VII Threats to Validity
We disabled compile-time optimizations in both Rust and LLVM to avoid missing bugs. However, we did not evaluate whether optimization would have had any effect on our results, and we did not determine whether any of the bugs we found had an effect on native execution.
@@
We did not use the same hardware when investigating the source of each bug as we did for large-scale data collection. Our value translation layer cannot detect when a binding uses an incorrect type that happens to have the correct layout, which may have caused us to miss bugs.
@@
Limited support for certain features, such as dynamically linked libraries, prevented us from achieving high test coverage; % of tests terminated due to an unsupported operation and % timed out. Additionally, the majority of libraries we examined linked against C, and not C++. However, these limitations did not prevent us from answering our research questions. Our goal was not to create a production-ready tool, but to understand the types of errors that can occur in Rust libraries that call foreign functions. The types of errors found in general may differ in unexpected ways from those found with our experimental setup.
@@ VIII Conclusion
We conducted a large-scale evaluation of Rust libraries to determine whether dynamic validation of Rust’s aliasing model is useful across foreign function boundaries. We created MiriLLI---a dynamic analysis tool that combines two existing interpreters---and we used it to identify new bugs from libraries. Though many of these bugs were aliasing violations, Rust’s Tree Borrows aliasing model was more permissive in this context than the earlier Stacked Borrows model. To ensure that these errors are easy to detect, the Rust project should invest in new dynamic analysis tools that can accommodate multi-language applications.
@@ Acknowledgment
We would like to thank Tomas Dougan for contributing to the implementation of MiriLLI. This material is based upon work supported by the Department of Defense and the National Science Foundation under Grant Nos. CCF-1901033, DGE1745016, and DGE2140739. Results presented in this paper were obtained using CloudBank [30], which is supported by the National Science Foundation under Grant No. CNS-1925001. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the Department of Defense or the National Science Foundation.

### v5

@@ IV-B Typing
We found 11 bugs in 10 crates involving typing or initialization. Most were caused by incorrect use of MaybeUninit<T>, but several involved incorrect foreign function bindings.
@@
We found four bugs that involved partial initialization of MaybeUninit<T>. Bugs #0 and #0 were found in the same library. In each case, an uninitialized instance of a struct T containing an array of values was created using MaybeUninit<T> and passed across the FFI to be initialized. The foreign call only initialized the first element of the array to zero. The remaining uninitialized elements would never be read by C, since every iteration stopped at the null terminator. Miri reported these initialization patterns as an error when MaybeUninit<T>::assume_init() was called. We fixed each bug by zero-initializing the entire array in Rust.
Bug #0 seemed to involve an equivalent pattern; MaybeUninit<T>::assume_init() triggered an error for a struct that was initialized by a foreign function. However, the function properly zero-initialized the struct using memset. After examining the LLVM bytecode of the foreign function, we found that all 88 bytes of the struct were initialized on the LLVM stack, but only the first 80 bytes were copied back into Rust. These missing bytes corresponded to padding fields, which were never directly accessed by the foreign library. Optimizations had been disabled, but LLVM still removed this unnecessary write access. We fixed Bug #0 by zero-initializing the padding fields before calling assume_init().
Incorrect Bindings
We found six crates with one or more incorrect foreign function bindings. All of these incorrect bindings had been written manually. Three of these crates had bindings with missing return types. In Bug #0, a binding was declared in a unit test without its 32-bit integer return type. The C implementation used this to return a status code indicating if one of its integer parameters was within bounds. The remaining two bugs involved bindings with incorrect integer types, which led to incomplete initialization. In Bug #0, a function was declared to return a 32-bit integer, but its C implementation returned a boolean value. In Bug #0, the last parameter of a function was declared as a 32-bit integer, but the C implementation expected size_t. The width of size_t is architecture-dependent, so this binding was only correct for 32-bit architectures. To fix this bug, multiple prior releases needed to be patched to use Rust’s usize type, which has similar semantics.
@@ IV-C Allocation
We found 12 issues related to allocation in 12 distinct crates. These include 10 memory leaks and one new cross-language deallocation bug. We also found one null pointer dereference, but it had no connection to Rust.
@@
Rust’s Box<T> and CString use similar APIs to encapsulate heap allocations. For each type, the function into_raw consumes an instance and produces a pointer to its interior heap allocation. To avoid a memory leak, this pointer must be reconstituted in its wrapper type using the function from_raw. We found four leaks caused by calling into_raw on either Box<T> or CString without later calling from_raw. Each type was used to allocate memory for a foreign function call, and each bug was fixed by adding a call to from_raw after the function returned. The remaining six leaks were of memory that had been allocated by C. Most were caused by neglecting to call a destructor that had been exposed by the C API. However, in Bug #0, the C API did not expose a destructor, since it had been designed based on the assumption that users would be able to call free. We could not use Rust’s std::alloc::dealloc to fix this, since that could lead to invalid cross-language deallocation.
@@ V Discussion
Our findings indicate that it is easy to inadvertently introduce undefined behavior in Rust libraries that call foreign functions. The errors we found can be prevented with careful auditing, but it seems likely that many will persist until the Rust community develops a production-ready method for finding aliasing violations in multi-language applications. We provide the following recommendations:
@@
The Rust community is in dire need of a production-ready solution for finding aliasing bugs in multi-language applications. Although our approach was capable of finding real-world bugs, its scalability is limited due to the requirement of implementing shims for system calls and the lack of a formal specification for the ABIs implemented by Rust, Clang, and LLVM. Instrumenting a shared, intermediate format is likely to be the most effective approach. The Krabcake project [33] (announced concurrently with MiriLLI) is developing an extension to Valgrind [22] that will provide support for detecting aliasing violations, but it is not yet capable of replicating our results. The Rust community should invest resources into completing a prototype implementation to evaluate in real-world applications. However, Valgrind can also incur significant runtime overhead. Our future work will involve creating an LLVM sanitizer that can detect aliasing violations, providing the performance necessary to find these errors at scale with fuzzing tools. We also expect that sophisticated static analysis tools would help developers avoid undefined behavior at foreign boundaries by issuing warnings for functions that copy pointers or mutate through pointers derived from immutable references.
@@ VI Related Work
Studies that examine how unsafe code is used in Rust libraries have consistently found that foreign function calls are a significant use case for unsafe code [4, 5]. In particular, Evans et al. [4] surveyed all published crates in September of 2018 and found that 22.5% of all unsafe function calls were to foreign functions. Rust developers also view foreign function calls as a central use case for unsafe code. Fulton et al. [6] interviewed 16 Rust developers and surveyed 178 developers, finding that nearly half of their participants used foreign function calls. Both Höltervennhoff et al. [7] and McCormack et al. [8] focused specifically on Rust developers who use unsafe code, and each found that the majority of participants used Rust’s FFI.
@@
Prior studies have examined open source contributions to categorize errors caused by unsafe code. Most considered foreign function calls to be out of scope, and none described aliasing violations in terms of Stacked or Tree Borrows. Both Qin et al. [34] and Xu et al. [35] examined bug and vulnerability reports from open source projects and found examples of allocation and typing errors that are similar to the ones that we discovered. Xu et al. [35] did encounter errors related to foreign function calls, but they were primarily ‘‘straightforward’’ issues related to application-specific invariants, layout, and alignment. Cui et al. [36] evaluated a taxonomy of 19 safety properties required by unsafe functions, but they made it an explicit design goal to exclude foreign functions.
@@
Prior approaches to static analysis [37, 34, 38, 39] have found a wide variety of bugs caused by improper use of unsafe code, but few approaches scale beyond Rust. Li et al. [40] applied dataflow analysis on LLVM IR and detected several memory leaks and cross-language deallocation errors in Rust libraries. We replicated several of the bugs they encountered. Hu et al. [41] modified several existing Rust analysis tools, including Miri, to analyze multi-language programs defined in a custom intermediate representation. However, they did not consider aliasing violations. Lei et al. [42] took a similar approach using WebAssembly as the target language, but their evaluation did not use Miri or consider aliasing issues, either. Recent deductive verifiers based on the RustBelt [43] model have used compositional verification [44] and automated proof search [45] to verify unsafe code, but none of these approaches reason about Stacked or Tree Borrows or have explicit support for foreign function calls.
@@ VII Threats to Validity
We disabled compile-time optimizations in both Rust and LLVM to avoid missing bugs. However, we did not evaluate whether optimization would have had any effect on our results, and we did not determine whether any of the bugs that we found had an effect on native execution.
@@
We did not use the same hardware when investigating the source of each bug as we did for large-scale data collection. Our value translation layer cannot detect when a binding uses an incorrect type that happens to have the correct layout, which may have caused us to miss bugs. Errors in Miri or other components of Rust’s nightly toolchain may have affected the correctness of our results. Since we manually investigated each bug, we have no reason to believe that these errors and their fixes are specific to any particular version of Rust’s toolchain, with the exception of Tree Borrows violations. The model was still evolving during our evaluation.
@@
Crates that could not compile with our nightly toolchain were excluded from our evaluation, even if they would have otherwise been compatible with MiriLLI. Limited support for certain features, such as dynamically linked libraries, also prevented us from achieving high test coverage; 61% of tests terminated due to an unsupported operation and 10% timed out. Additionally, the majority of libraries we examined linked against C, and not C++. Crates that failed to compile using our nightly toolchain were excluded from the evaluation, as well as crates that overrode the system’s default C and C++ compiler. However, these limitations did not prevent us from answering our research questions. Our goal was not to create a production-ready tool, but to understand the types of errors that can occur in Rust libraries that call foreign functions. The bugs that occur in general may differ in unexpected ways from those found with our experimental setup.
@@ VIII Conclusion
We conducted a large-scale evaluation to investigate the types of undefined behavior that occur in Rust libraries that call foreign functions. We created MiriLLI---a dynamic analysis tool that combines two existing interpreters---and we used it to identify 47 new bugs from 37 libraries. Though many of these bugs were aliasing violations, Rust’s Tree Borrows aliasing model was more permissive in this context than the earlier Stacked Borrows model. To ensure that these errors are easy to detect, the Rust project should invest in new dynamic analysis tools that can accommodate multi-language applications.
@@ Acknowledgment
We thank Tomas Dougan for contributing to the implementation of MiriLLI. This material is based on work supported by the Department of Defense and the National Science Foundation under Grant Nos. CCF-1901033, DGE1745016, and DGE2140739. Results presented in this paper were obtained using CloudBank [30], which is supported by the National Science Foundation under Grant No. CNS-1925001. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the Department of Defense or the National Science Foundation.